OpenStudy (pina_colada123):

Solve the equation for y. 8x – 9y = 11 A) y= -9/8x-9/11 B) y= 9/8x+11/8 C) y= -8x-11 D) y=8/9x-11/9

rishavraj (rishavraj):

first thing u need to do is isolate 9y....so subtract 8x on both sides .... so it would be -9y = 11 - 8x then multiply by -1 on both sides to get rid of -ve sign with y

rishavraj (rishavraj):

@pina_colada123 ???

OpenStudy (pina_colada123):

agg confused right now i have -9y=11-8x

OpenStudy (pina_colada123):

what do i do after @rishavraj

rishavraj (rishavraj):

yup -9y = 11 - 8x so 9y = 8x - 11 u got tht???@pina_colada123

rishavraj (rishavraj):

@pina_colada123

OpenStudy (pina_colada123):

yes

rishavraj (rishavraj):

now u need to divide by 9 on both sides....to isolate y

rishavraj (rishavraj):

there????? @pina_colada123

OpenStudy (pina_colada123):

ya so it would be D

rishavraj (rishavraj):

yup :))) @pina_colada123

OpenStudy (pina_colada123):

ya thanx